OpenStudy (anonymous):

find the horizontal tangeant of y=4x^2+16x+21

hartnn (hartnn):

could you find dy/dx ?

OpenStudy (anonymous):

I figured it out!! :D is it x= -2?

hartnn (hartnn):

thats the x-coordinate of the horizonatal tangent the equation of horizontal tangent will be of the form y= constant so, from x=2 and your original equation, y=... you just need to find the y co-ordinate

OpenStudy (anonymous):

could you help me on y=x^4-2x^2+1 instead? (:

OpenStudy (anonymous):

so just plug it in? :)

hartnn (hartnn):

yes, for the 1st problem just put x=-2 in your equation y=4x^2+16x+21

hartnn (hartnn):

for y=x^4-2x^2+1 what did u get dy/dx as ?

OpenStudy (anonymous):

4x^3-4x

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

did u get y=5 as your final answer for 1st ? yes, thats correct 4x^3-4x =0 solve this for x

OpenStudy (anonymous):

using quadratic?

OpenStudy (anonymous):

or how? lol

hartnn (hartnn):

factor out the 'x' first 4x (x^2-1) = 0 x=0 , or x^2-1 =0

OpenStudy (anonymous):

oohhh! is it x=0, x=1, x=-1 ?

hartnn (hartnn):

sorry for late replies yes, those are correct find y value from that

OpenStudy (anonymous):

okay, thanks(: